# The number of ends of big mapping class groups

### v1

@@ Abstract.
Big mapping class groups, number ends of groups
@@ 1. Introduction
A central theme in geometric group theory is the large-scale geometry of groups. By this we mean the metric structure that is preserved under quasi-isometries—maps which preserve distances up to a uniformly bounded multiplicative and additive error (Section 2.1). In a 1983 ICM address [gromov-icm], Gromov set forth a far-reaching program advocating the study of finitely generated groups as geometric objects and their classification up to quasi-isometry. Among the earliest and most classical quasi-isometric invariants is the number of ends. Informally speaking, the number of ends of a simplicial graph is the number of topologically distinct ways to escape to infinity, and the number of ends of a finitely generated group is defined to be the number of ends of its Cayley graph with respect to some (any) finite generating set (Section 2.2). Freudenthal [Freudenthal] and Hopf [Hopf] independently proved that the number of ends of a finitely generated infinite group is always either one, two, or infinite, where the group has exactly two ends if and only if it contains an infinite cyclic subgroup of finite index. A celebrated theorem of Stallings [Stallings68, MR415622] provides a complete characterization of finitely generated groups with infinitely many ends. It states that a finitely generated group has more than one end if and only if it splits over finite groups, or equivalently, it admits an action on a simplicial tree with finite edge-stabilizers and without edge-inversions or a global fixed vertex. More generally, infinite countable groups, even when not finitely generated, enjoy the same trichotomy on their number of ends [yves].
@@
It is a classical fact that the mapping class groups of finite-type surfaces, except for some cases when the surface has low complexity, are all one-ended (many even have geodesic divergence bounded above by a quadratic function [DuchinRafi]). So the goal of this paper is to investigate the number of ends of big mapping class groups, i.e., mapping class groups of infinite-type surfaces (Section 2.3). Interest in big mapping class groups grew significantly after Bavard [Bavard16] proved that the ray graph is Gromov hyperbolic, analogous to the famous theorem of Masur–Minsky [MasurMinsky] that the curve graph is Gromov hyperbolic. Since then, much effort has been made to understand big mapping class groups and to compare them with the better-studied mapping class groups of finite-type surfaces (see for example, [Long, Wu1, Israel]).
@@
Geometric group theory typically concerns finitely (or compactly) generated groups because it is well-known that the word metrics associated to any two finite (or compact) generating sets are quasi-isometric to each other. Big mapping class groups are not finitely or countably infinitely generated, nor are they compactly generated. So a priori, they do not have a well-defined geometry. However, Rosendal [Rosendal] developed ideas which allow for the study of the large-scale geometry of topological groups which are not locally compact, let alone finitely generated. Specifically, Rosendal proved that for groups which admit a coarsely bounded (CB) generating set (Section 2.5), it is also the case that the word metrics associated to any two CB generating sets are quasi-isometric to each other. We remark that a subset of a locally compact space is CB if and only if it is compact, so this notion does indeed generalize older ones. Recent work of Mann–Rafi [MannRafi] applied this idea to classify (many of) the big mapping class groups which are CB generated, and therefore have a well-defined quasi-isometry type. Using their tools and classification, several results have shown that these groups admit a rich geometric structure ([GRV, HQR]).
@@
Within this framework, one can now reasonably define the number of ends of a CB generated group to be the number of ends of its Cayley graph with respect to some (any) CB generating set. Then it is natural to ask: How many ends do CB generated big mapping class groups have? A theorem in [OhPengitore22] states that the number of ends of a connected, unbounded, (coarsely) transitive graph is one, two, or infinite, generalizing the theorem of Freudenthal–Hopf. The theorem statement includes an assumption that the graph is locally finite, but in fact, this assumption is unnecessary as it is never used in the proof. The Cayley graph of a CB generated mapping class group is connected and transitive. So even though it is locally infinite (in general, the degree of each vertex is uncountable), we may conclude the following.
@@
A surface $\Sigma$ of infinite type falls into one of two classes depending on whether or not it contains a finite-type non-displaceable subsurface—a subsurface which intersects its image under every homeomorphism of $\Sigma$. According to recent work, the large-scale geometry of the big mapping class group $\operatorname{Map}(\Sigma)$ often depends on whether or not $\Sigma$ contains a non-displaceable subsurface of finite type. For example, if $\Sigma$ does contain a non-displaceable subsurface of finite type, then $\operatorname{Map}(\Sigma)$ is not coarsely bounded [MannRafi] and it admits a continuous and nonelementary action by isometries on a Gromov hyperbolic space [HQR] (the authors also prove that under an extra condition, $\operatorname{Map}(\Sigma)$ admits such an action only if $\Sigma$ contains a non-displaceable subsurface). On the other hand, if $\Sigma$ does not contain a non-displaceable subsurface and $\operatorname{Map}(\Sigma)$ is not CB, then $\operatorname{Map}(\Sigma)$ has infinite asymptotic dimension [GRV].
In an effort to further classify the number of ends of big mapping class groups, we focus on those surfaces $\Sigma$ which do not contain a non-displaceable subsurface. Since we are attempting to analyze the large-scale geometry of $\operatorname{Map}(\Sigma)$, we are only interested in the surfaces for which $\operatorname{Map}(\Sigma)$ is CB generated, but not CB itself. A broad class of surfaces which satisfy these conditions is the class of avenue surfaces (Section 2.6). Avenue surfaces were defined in [HQR], and the only additional condition imposed on them is a weak topological condition called tameness on the space of ends of $\Sigma$. Tameness is rather technical and lengthy to define, so we choose to work within the slightly more restrictive, but much more easily defined, class of stable surfaces. Stable surfaces form a substantial subclass of tame surfaces, and it has become somewhat standard to work with them. For example, several other works in the literature [Bar-NatanVerberne, GRV, BDR] similarly restrict to stable surfaces in order to conveniently work within the framework developed in [MannRafi]. That being said, we expect our results to also hold in the setting of surfaces with tame end spaces.
Every avenue surface has two distinct ends $e_{+}$ and $e_{-}$ which are maximal with respect to the partial pre-order $\preceq$ on ${\mathrm{End}}(\Sigma)$ defined in [MannRafi]. It follows from [Schaffer-Cohen] that for every end $x$ which is an immediate predecessor of $e_{+}$ (and $e_{-}$), the equivalence class of $x$ is either a discrete set or a totally disconnected perfect set, where the equivalence relation is defined by $\preceq$. In the first case, we say the end $x$ is of discrete type. We can now state our main theorem.
@@
If $\Sigma$ is a stable avenue surface with at least one discrete-type end, then $\operatorname{Map}(\Sigma)$ is one-ended.
The most basic example of a surface covered by this theorem is the bi-infinite flute surface. This surface indeed has exactly two maximal ends, and all other ends, represented by punctures, form a single equivalence class homeomorphic to $\mathbb{Z}$. On the other hand, one surface which is not covered by our theorem is the Jacob’s ladder surface, which has exactly two ends, each accumulated by genus. It may seem at first that the Jacob’s ladder surface should be no more difficult to handle than the bi-infinite flute surface. However, one significant difference is that in the bi-infinite flute surface, curves cannot meaningfully “interact” with the punctures, since the punctures are not points on the surface, but in the Jacob’s ladder surface, curves may “interact” with the handles in a highly non-trivial way. We do not know whether our methods may be generalized to deal with this additional layer of complexity, so we ask the following.
@@
A common strategy to study mapping class groups is to construct appropriate metric spaces on which the groups act with good geometric properties. For finite-type surfaces, these spaces are often graphs whose vertices are represented by curves and arcs on the surface, and the fundamental example is the curve graph defined in [Harvey]. Avenue surfaces fall in the broader class of translatable surfaces, for which Schaffer-Cohen [Schaffer-Cohen] introduced the translatable curve graph, whose vertices correspond to simple closed curves which separate the two maximal ends, and whose edges connect pairs of vertices whenever the corresponding curves bound a subsurface homeomorphic to one of a pre-chosen finite collection of subsurfaces (Section 2.7). The translatable curve graph is not only an analogue of the curve graph in the setting of translatable surfaces; it is even quasi-isometric to the associated mapping class group [Schaffer-Cohen]. So every property which is invariant under quasi-isometry holds for the mapping class group precisely when it holds for the translatable curve graph. Therefore, we directly analyze the translatable curve graph in order to determine the number of ends of the mapping class group. To this end, we construct an operation called lassoing that takes a vertex in the graph to an adjacent vertex and, more generally, takes a (well-behaved) path to an adjacent path. Furthermore, we find some lower bounds, which we call flux and Hamming distance, on the graph distance which are crucial to guaranteeing that points are sufficiently far from a given basepoint. Section 3 presents a proof of the theorem in the case of the bi-infinite flute, capturing many of the key ideas, and the theorem is proved in full in Section 4.
We finish with some more questions that arose during our investigation.
@@
In particular, this question is open for the specific surface studied in [Bavard16] that initiated much of the recent interest in big mapping class groups.
@@ 2.3. Surfaces and the mapping class group
A surface is a connected, orientable, $2$-dimensional topological manifold without boundary. A surface $\Sigma$ is of finite type if its fundamental group is finitely generated; otherwise $\Sigma$ is of infinite type. A subsurface $S$ of a surface $\Sigma$ is a closed, connected subset of $\Sigma$ which is itself a surface with boundary. Assume the boundary of a subsurface consists of a finite number of pairwise disjoint simple closed curves, and that none of these boundary curves bounds a disk or a punctured disk in $\Sigma$.
Denote by $\operatorname{Homeo}^{+}(\Sigma)$ the group of orientation-preserving homeomorphisms of $\Sigma$ with the compact-open topology, and denote by $\operatorname{Homeo}_{0}(\Sigma)$ the connected component of the identity in $\operatorname{Homeo}^{+}(\Sigma)$. The mapping class group of $\Sigma$ is defined as the group $\operatorname{Map}(\Sigma)=\operatorname{Homeo}^{+}(\Sigma)/\operatorname{Homeo}_{0}(\Sigma)$ of all isotopy classes of orientation-preserving homeomorphisms of $\Sigma$, and $\operatorname{Map}(\Sigma)$ is endowed with the quotient topology. When $\Sigma$ is of infinite type, $\operatorname{Map}(\Sigma)$ is a Polish group which is not locally compact.
@@ 2.4. The space of ends of a surface
The space of ends ${\mathrm{End}}(\Sigma)$ of a surface $\Sigma$ is the inverse limit of the system of components of complements of compact subsets of $\Sigma$. Intuitively, each end corresponds to a way of leaving every compact subset of $\Sigma$ (see [Richards] for details). Denote by ${\mathrm{End}}^{g}(\Sigma)\subseteq{\mathrm{End}}(\Sigma)$ the subspace of ends which are accumulated by genus, and denote by $\operatorname{genus}(\Sigma)$ the (possibly infinite) genus of $\Sigma$. By a theorem of Richards [Richards], connected, orientable surfaces $\Sigma$ are classified up to homeomorphism by the triple $(\operatorname{genus}(\Sigma),{\mathrm{End}}(\Sigma),{\mathrm{End}}^{g}(\Sigma))$. For a subsurface $S\subseteq\Sigma$, the space of ends of $S$ is defined similarly and is denoted by ${\mathrm{End}}(S)$. The embedding of $S$ in $\Sigma$ gives a natural embedding of ${\mathrm{End}}(S)$ into ${\mathrm{End}}(\Sigma)$.
Assume that $\Sigma$ is a surface of infinite type. One of the key tools used in [MannRafi] was a partial pre-order $\preceq$ on ${\mathrm{End}}(\Sigma)$. By definition, $y\preceq x$ if, for every clopen neighborhood $U$ of $x$, there exists a clopen neighborhood $V$ of $y$ which is homeomorphic to a subset of $U$. This induces an equivalence relation on ${\mathrm{End}}(\Sigma)$, where $x$ and $y$ are equivalent ends if $x\preceq y$ and $y\preceq x$, and we denote by $E(x)$ the equivalence class of $x$. Denote by $\mathcal{M}({\mathrm{End}}(\Sigma))$ the subset of ${\mathrm{End}}(\Sigma)$ of ends which are maximal with respect to $\preceq$, that is, the set of ends $x\in{\mathrm{End}}(\Sigma)$ such that $x\preceq y$ implies $y\in E(x)$. The elements of $\mathcal{M}({\mathrm{End}}(\Sigma))$ are called maximal ends.
For an end $x\in{\mathrm{End}}(\Sigma)$, a clopen neighborhood $U$ of $x$ is stable if for any smaller neighborhood $U^{\prime}\subset U$ of $x$, there is a homeomorphic copy of $U$ contained in $U^{\prime}$. (see [BDR, Proposition 3.2] for equivalent definitions of a stable neighborhood). The surface $\Sigma$ is stable if every end $x\in{\mathrm{End}}(\Sigma)$ has a stable neighborhood.
@@ 2.5. Geometry of big mapping class groups
Let $G$ be a CB generated Polish group, and fix a symmetric CB generating set $S$. The number of ends of $G$ is the number of ends of the Cayley graph of $G$ with respect to $S$.

### v2

@@ Abstract.
Big mapping class groups, number of ends of groups
@@ 1. Introduction
A central theme in geometric group theory is the large-scale geometry of groups. By this we mean the metric structure that is preserved under quasi-isometries—maps which preserve distances up to a uniformly bounded multiplicative and additive error (Section 2.1). In a 1983 ICM address [gromov-icm], Gromov advocated the study of finitely generated groups as geometric objects and their classification up to quasi-isometry. Among the earliest and most classical quasi-isometric invariants is the number of ends. The number of ends of a simplicial graph is the number of topologically distinct ways to escape to infinity, and the number of ends of a finitely generated group is defined to be the number of ends of its Cayley graph with respect to some (any) finite generating set (Section 2.2). Freudenthal [Freudenthal] and Hopf [Hopf] independently proved that the number of ends of a finitely generated infinite group is always either one, two, or infinite, and the group has exactly two ends if and only if it is virtually infinite cyclic. A celebrated theorem of Stallings [Stallings68, MR415622] provides a complete characterization of the case of infinitely many ends. It states that a finitely generated group has more than one end if and only if it splits over a finite group, or equivalently, it admits an action on a simplicial tree with finite edge-stabilizers and without edge-inversions or a global fixed vertex. More generally, countably infinite groups, even when not finitely generated, enjoy the same trichotomy on their number of ends [yves].
@@
It is a classical fact that mapping class groups of finite-type surfaces, except for some cases when the surface has low complexity, are all one-ended (all but a few even have geodesic divergence bounded above by a quadratic function [DuchinRafi]). So the goal of this paper is to investigate the number of ends of big mapping class groups, i.e., mapping class groups of infinite-type surfaces (Section 2.3). Interest in big mapping class groups grew significantly after Bavard [Bavard16] proved that the ray graph is Gromov hyperbolic, analogous to the famous theorem of Masur–Minsky [MasurMinsky] that the curve graph is Gromov hyperbolic. Since then, much effort has been made to understand big mapping class groups and to compare them with the better-studied mapping class groups of finite-type surfaces (see for example, [Long, Wu1, Israel]).
@@
Geometric group theory typically concerns finitely (or compactly) generated groups because it is well-known that the word metrics associated to any two finite (or compact) generating sets are quasi-isometric to each other. Big mapping class groups are not finitely or countably infinitely generated, nor are they compactly generated. So a priori, they do not have a well-defined geometry. However, Rosendal [Rosendal] developed ideas which allow for the study of the large-scale geometry of topological groups which may not be locally compact, let alone finitely generated. Specifically, Rosendal proved that for groups which admit a coarsely bounded (CB) generating set (Section 2.5), it is also the case that the word metrics associated to any two CB generating sets are quasi-isometric to each other. We remark that a subset of a locally compact space is CB if and only if it is compact, so this notion does indeed generalize older ones. Recent work of Mann–Rafi [MannRafi] applied this idea to classify (many of) the big mapping class groups which are CB generated, and therefore have a well-defined quasi-isometry type. Using their tools and classification, several results have shown that these groups admit a rich geometric structure [GRV, HQR, RoMo25]. See also [Hill25] for the case of pure mapping class groups.
@@
Within this framework, the number of ends of a CB generated group may be defined to be the number of ends of its Cayley graph with respect to some (any) CB generating set. Then it is natural to ask: How many ends do CB generated big mapping class groups have? A theorem in [OhPengitore22] states that the number of ends of a connected, unbounded, (coarsely) transitive graph is one, two, or infinite, generalizing the theorem of Freudenthal–Hopf. The theorem statement includes an assumption that the graph is locally finite, but in fact, this assumption is unnecessary as it is never used in the proof. The Cayley graph of a CB generated mapping class group is connected and transitive. So even though it is locally infinite (in general, the degree of each vertex is uncountable), we may conclude the following.
@@
A surface $\Sigma$ of infinite type falls into one of two classes depending on whether or not it contains a finite-type non-displaceable subsurface—a subsurface which intersects its image under every homeomorphism of $\Sigma$. According to recent work, the large-scale geometry of the big mapping class group $\operatorname{Map}(\Sigma)$ often depends on whether or not $\Sigma$ contains a non-displaceable subsurface of finite type. For example, if $\Sigma$ does contain a non-displaceable subsurface of finite type, then $\operatorname{Map}(\Sigma)$ is not CB [MannRafi] and it admits a continuous and nonelementary action by isometries on a Gromov hyperbolic space [HQR] (the authors also prove that under an extra condition, $\operatorname{Map}(\Sigma)$ admits such an action only if $\Sigma$ contains a non-displaceable subsurface). On the other hand, if $\Sigma$ does not contain a non-displaceable subsurface of finite type and $\operatorname{Map}(\Sigma)$ is not CB, then $\operatorname{Map}(\Sigma)$ has infinite asymptotic dimension [GRV].
In an effort to further classify the number of ends of big mapping class groups, we focus on those surfaces $\Sigma$ which do not contain a non-displaceable subsurface of finite type. Since we are attempting to analyze the large-scale geometry of $\operatorname{Map}(\Sigma)$, we are only interested in the surfaces for which $\operatorname{Map}(\Sigma)$ is CB generated, but not CB itself. A broad class of surfaces which satisfy these conditions is the class of avenue surfaces (Section 2.6). Avenue surfaces were defined in [HQR], and the only additional condition imposed on them is a weak topological condition called tameness on the space of ends of $\Sigma$. Tameness is rather technical and lengthy to define, so we choose to work within the slightly more restrictive, but much more easily defined, class of stable surfaces (Section 2.4). Stable surfaces form a substantial subclass of tame surfaces, and it has become somewhat standard to work with them. For example, several other works in the literature [Bar-NatanVerberne, GRV, BDR] similarly restrict to stable surfaces in order to conveniently work within the framework developed in [MannRafi]. That being said, we expect our results to also hold for surfaces with tame end space.
Every avenue surface has two distinct ends $e_{+}$ and $e_{-}$ which are maximal with respect to the partial pre-order $\preceq$ on ${\mathrm{End}}(\Sigma)$ defined in [MannRafi]. It follows from [Schaffer-Cohen] that for each end $x$ which is an immediate predecessor of $e_{+}$ (and $e_{-}$), the equivalence class of $x$ is either a discrete set or a totally disconnected perfect set, where the equivalence relation is defined by $\preceq$. In the first case, we say the end $x$ is of discrete type. We can now state our main theorem.
@@
If $\Sigma$ is a stable avenue surface with a discrete-type end, then $\operatorname{Map}(\Sigma)$ is one-ended.
The most basic example of a surface covered by this theorem is the bi-infinite flute surface. This surface indeed has exactly two maximal ends, and all other ends, represented by punctures, form a single equivalence class homeomorphic to $\mathbb{Z}$. On the other hand, one surface which is not covered by our theorem is the Jacob’s ladder surface, which has exactly two ends, each accumulated by genus. It may seem at first that the Jacob’s ladder surface should be no more difficult to handle than the bi-infinite flute surface. However, one significant difference is that in the bi-infinite flute surface, curves cannot meaningfully “interact” with the punctures, since the punctures are not points on the surface, but in the Jacob’s ladder surface, curves may “interact” with the handles in a highly non-trivial way. We do not know whether our methods may be modified to deal with this additional layer of complexity, so we ask the following.
@@
A common strategy to study mapping class groups is to construct appropriate metric spaces on which the groups act with good geometric properties. For finite-type surfaces, these spaces are often graphs whose vertices are represented by curves and arcs on the surface, and the fundamental example is the curve graph defined in [Harvey]. Avenue surfaces fall in the broader class of translatable surfaces, for which Schaffer-Cohen [Schaffer-Cohen] introduced the translatable curve graph, whose vertices correspond to simple closed curves which separate the two maximal ends, and whose edges connect pairs of vertices whenever the corresponding curves bound a subsurface homeomorphic to one of a pre-chosen finite collection of subsurfaces (Section 2.7). The translatable curve graph is not only an analogue of the curve graph in the setting of translatable surfaces; it is even quasi-isometric to the associated mapping class group [Schaffer-Cohen]. So every property which is invariant under quasi-isometry holds for the mapping class group precisely when it holds for the translatable curve graph. Therefore, we directly analyze this graph in order to determine the number of ends of the mapping class group. To this end, we devise an operation called lassoing (Section 4.1) that takes a vertex in the graph to an adjacent vertex and, more generally, takes a (well-behaved) path to an adjacent path. Furthermore, we find some lower bounds on the graph distance, called flux (Section 4.1) and Hamming distance (Section 4.3), that are crucial to proving the graph is one-ended. Section 3 presents a proof of the theorem in the case of the bi-infinite flute, capturing many of the key ideas, and the theorem is proved in full in Section 4.
We finish with additional questions that arose during our investigation.
@@
In particular, this question is open for the particular surface studied in [Bavard16] that initiated much of the recent interest in big mapping class groups.
@@ 2.3. Surfaces and the mapping class group
A surface $\Sigma$ is a connected, orientable, $2$-dimensional topological manifold without boundary. It is of finite type if its fundamental group is finitely generated; otherwise it is of infinite type. A subsurface $S$ of $\Sigma$ is a closed, connected subset of $\Sigma$ which is itself a surface with boundary. Assume the boundary of a subsurface consists of a finite number of pairwise disjoint simple closed curves, none of which bound a disk or a punctured disk in $\Sigma$.
Denote by $\operatorname{Homeo}^{+}(\Sigma)$ the group of orientation-preserving homeomorphisms of $\Sigma$, with the compact-open topology, and denote by $\operatorname{Homeo}_{0}(\Sigma)$ the connected component of the identity in $\operatorname{Homeo}^{+}(\Sigma)$. The mapping class group of $\Sigma$ is defined as the group $\operatorname{Map}(\Sigma)=\operatorname{Homeo}^{+}(\Sigma)/\operatorname{Homeo}_{0}(\Sigma)$ of all isotopy classes of orientation-preserving homeomorphisms of $\Sigma$, and $\operatorname{Map}(\Sigma)$ is endowed with the quotient topology. When $\Sigma$ is of infinite type, $\operatorname{Map}(\Sigma)$ is a Polish group which is not locally compact.
@@ 2.4. The space of ends of a surface
The space of ends ${\mathrm{End}}(\Sigma)$ of a surface $\Sigma$ is the inverse limit of the system of components of complements of compact subsets of $\Sigma$. Intuitively, each end corresponds to a way of leaving every compact subset of $\Sigma$ (see [Richards] for details). Denote by ${\mathrm{End}}^{g}(\Sigma)\subset{\mathrm{End}}(\Sigma)$ the subspace of ends which are accumulated by genus, and denote by $\operatorname{genus}(\Sigma)$ the (possibly infinite) genus of $\Sigma$. By a theorem of Richards [Richards], connected, orientable surfaces $\Sigma$ are classified up to homeomorphism by the triple $(\operatorname{genus}(\Sigma),{\mathrm{End}}(\Sigma),{\mathrm{End}}^{g}(\Sigma))$. For a subsurface $S\subset\Sigma$, the space of ends of $S$ is defined similarly and is denoted by ${\mathrm{End}}(S)$. The embedding of $S$ in $\Sigma$ gives a natural embedding of ${\mathrm{End}}(S)$ into ${\mathrm{End}}(\Sigma)$.
Assume that $\Sigma$ is a surface of infinite type. One of the key tools used in [MannRafi] was a partial pre-order $\preceq$ on ${\mathrm{End}}(\Sigma)$. By definition, $y\preceq x$ if, for every clopen neighborhood $U$ of $x$, there exists a clopen neighborhood $V$ of $y$ which is homeomorphic to a subset of $U$. This induces an equivalence relation on ${\mathrm{End}}(\Sigma)$, where $x$ and $y$ are equivalent ends if $x\preceq y$ and $y\preceq x$, and we denote by $E(x)$ the equivalence class of $x$. Denote by $\mathcal{M}({\mathrm{End}}(\Sigma))$ the set of ends of $\Sigma$ which are maximal with respect to $\preceq$, that is, the set of ends $x\in{\mathrm{End}}(\Sigma)$ such that $x\preceq y$ implies $y\in E(x)$. The elements of $\mathcal{M}({\mathrm{End}}(\Sigma))$ are called maximal ends.
For an end $x\in{\mathrm{End}}(\Sigma)$, a clopen neighborhood $U$ of $x$ is stable if for any smaller neighborhood $U^{\prime}\subset U$ of $x$, there is a homeomorphic copy of $U$ contained in $U^{\prime}$. (see [BDR, Proposition 3.2] for equivalent definitions of a stable neighborhood). The surface $\Sigma$ is stable if each of its ends has a stable neighborhood.
@@ 2.5. Geometry of big mapping class groups
Let $G$ be a CB generated Polish group, and fix a symmetric CB generating set $S$. The number of ends of $G$ is the number of ends of its Cayley graph with respect to $S$.